OpenStudy (anonymous):

Rearrange x=2/3 y squared +4

OpenStudy (anonymous):

ready to work on this together?

OpenStudy (anonymous):

that would be great

OpenStudy (anonymous):

Sweeet! Ok, first thing we want to do is to try to get the y on one side and the rest of the junk on the other side of the equals sign.

OpenStudy (anonymous):

Easiest thing to remove is the +5. How do we get rid of it from the right side of the equals sign?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

2/3y^2=x-4 multiply by 3/2 next?

OpenStudy (anonymous):

Oooohhhh I'm sorry. I mixed you up with someone else. -4 IS correct. Sorry!

OpenStudy (anonymous):

You're absolutely correct. multipy by 3/2 (which is technically the same as dividing by 2/3)

OpenStudy (anonymous):

y^2=3/2x-4 *then remove the root?

OpenStudy (anonymous):

power of two i mean

OpenStudy (anonymous):

Yes, to do this, you square root everything on both sides.

OpenStudy (anonymous):

y=sqrt of 3/2x-4 ????

OpenStudy (anonymous):

wolfram alpha gives me a different solution to that

OpenStudy (anonymous):

Weird. Lemme write this out in pencil then.

OpenStudy (anonymous):

I got the same answer. :/ Scroll down to where it says "answers for y"

OpenStudy (anonymous):

they give me this....is that the exact same sort of answer as what i got?

OpenStudy (anonymous):

Yup! They just put a separate square root on each expression. It's still the same answer.

OpenStudy (anonymous):

their answer actually appears to look like sqrt3/2 multiplied by sqrtx-4

OpenStudy (anonymous):

Basically, they didn't distribute the square root in wolfram alpha.

OpenStudy (anonymous):

Yup! Same thing :D

OpenStudy (anonymous):

For example: Square root of 4 times square root of four. How much is that?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

I had the right answer the first time I did it. I guess I just needed to confirm it!

OpenStudy (anonymous):

and if you combined the suare root and said "square root of (4 times 4)"?

OpenStudy (anonymous):

No problemo.

OpenStudy (anonymous):

I sorta assumed that, just good to hear someone else confirm it ie "sqrt(4 x 4)

OpenStudy (anonymous):

hey uber, I asked this question last night and "mathlearner" gave me a different answer....I think his is wrong. http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e180a500b8bc2275746b45b

OpenStudy (anonymous):

Hmmm.... it might be correct, but that is very weird notation. Lemme double check.

OpenStudy (anonymous):

According to wolfram, it's false. http://www.wolframalpha.com/input/?i=x%3D+2%2F3+y%5E2+%2B+4+%3D+sqrt%282-4x%2F3x%29

OpenStudy (anonymous):

I think he has too many X's. They should cancel out the way he wrote it.

OpenStudy (anonymous):

2x/3x should just equal 2/3 with no x

OpenStudy (anonymous):

you're right.he does have one too many Xs.